# Efficient Fairness-Performance Pareto Front Computation

### v1

@@ H.1 The Relation between Representation and Classifier Pareto Fronts
Let the uncertainy measure $h$ be defined by (4).
Lemma H.1.
@@
$\mathbb{E}_{z\sim Z}h(\mathbb{P}\left(Y|Z=z\right))\leq\epsilon(\hat{Y})\text{and }\left\|\mu_{0}-\mu_{1}\right\|_{TV}\leq\Delta_{SP}(\hat{Y}).$ (118)
@@
[…]
@@
We now turn to the first statement, 118. Let $\hat{Y}$ be a classifier of $Y$ as a function of $(X,A)$. Observe that thus by definition $\mathbb{P}\left(\hat{Y}|X,A\right)$ induces a distribution on the set $\left\{0,1\right\}$, and thus may be considered as a representation $Z:=\hat{Y}$ of $(X,A)$ on that set. We now relate the properties of this $Z$ as a representation to the quantities $\epsilon(\hat{Y})$ and $\Delta_{SP}(\hat{Y})$. Similarly to the argument above, the first part of (118) follows since $\mathbb{E}_{z\sim Z}h(\mathbb{P}\left(Y|Z=z\right))$ is the best possible error over all classifiers. For the second part, note that since $\hat{Y}$ is binary, we have
@@
H.2 Comparison to Common Fair Classifiers
In this Section we evaluated the accuracy-fairness tradeoff for common fair classification algorithms using three datasets - the Adult dataset (income prediction), COMPAS (recidivism prediction), and LSAC (law school admission). These datasets are standard fair classification benchmarks, and have relevance to real-world decision-making processes.
Fairness classifiers are generally categorized into three groups: fair pre-processing, fair in-processing, and fair post-processing. Fair pre-processing and post-processing utilize standard classifiers as part of the fair classification pipeline. In the experiments below we have evaluated two widely adopted algorithms, representing two different categories: (i) FairGBM Cruz et al., (2023), an in-processing method where a boosting trees algorithm (LightGBM) is subject to pre-defined fairness constraints, and (ii) Balanced-Group-Threshold, a post-processing method which adjusts the threshold per group to obtain a certain fairness criterion. For FairGBM we have used the original implementation provided by the authors, while for Balanced-Group-Threshold post-processing we have used the implementations available via Aequitas, a popular bias and fairness audit toolkit, Saleiro et al., (2018).
It is important to note that, as a rule, common fairness classification methods are not designed to control the fairness-accuracy trade-off explicitly. Instead, in most cases, these methods rely on rerunning the algorithm for a wide range of hyperparameter settings, in the hope that different hyperparameters would result in different fairness-accuracy trade-off points. However, there typically is no direct known and controlled relation between hyperpatameters and the obtained fairness-accuracy trade-off.
For FairGBM, we utilized the hyperparameter ranges specified in the original paper, Cruz et al., (2023). In the case of the balancing post-processing method, we iterated over all possible parameters to ensure a comprehensive analysis.
Figure 4 shows the MIFPO computed Pareto front, and all hyperparameter runs of the two algorithms above, with accuracy evaluated on the test set. These experiments demonstrate the following two points: (a) The standard classifiers achieve a considerably lower accuracy than what is theoretically possible at a given fairness level. (b) the existing methods are also unable to present solutions for the full range of the statistical parity values. The values from the FGBM and the post-processing algorithms all have statistical parity $\leq 0.2$.
The above results emphasize the limitations of current fair classification algorithms in achieving optimal trade-offs between accuracy and fairness across the full range of fairness values, and suggest that there is room for improvement in developing more flexible and comprehensive fair classification techniques.

### v2

@@ Appendix I Fair Classifiers As Fair Representations
Let the uncertainy measure $h$ be defined by (4). Note that the first part of the Lemma does use the special properties of this $h$ and does not necessarily hold for other costs $h$.
Lemma I.1.
@@
[…]
@@
Let us begin with the second part of the Lemma, inequalities (109). Given a representation $Z$, $\epsilon(\hat{Y})\leq\mathbb{E}_{z\sim Z}h(\mathbb{P}\left(Y|Z=z\right))$ follows since $\mathbb{E}_{z\sim Z}h(\mathbb{P}\left(Y|Z=z\right))$ is the error of the optimal classifier of $Y$ as a function of $Z$. We choose $\hat{Y}$ to be such an optimal classifier and thus satisfy the above inequality, with equality. Next, the second inequality in (109) holds for for any classifier $\hat{Y}$ derived from $Z$. The argument below is a slight generalisation of the argument in Madras et al., (2018). Define $f(z)=\mathbb{P}\left(\hat{Y}=1|Z=z\right)$. Note that for $a\in\left\{0,1\right\}$, $\mathbb{P}\left(\hat{Y}=1|A=a\right)=\int f(z)d\mu_{a}(z)$. Thus
$\mathbb{P}\left(\hat{Y}=1|A=0\right)-\mathbb{P}\left(\hat{Y}=1|A=1\right)$ $=\int f(z)d\mu_{0}(z)-\int f(z)d\mu_{1}(z)$ (110)
$\leq\sup_{g|\left|g\right|\leq 1}\left|\int g(z)d\mu_{0}(z)-\int g(z)d\mu_{1}(z)\right|$ (111)
@@
[…]
@@
Appendix J Computational Complexities
In this Section we discuss alternative discretization schemes for the MIFPO algorithm. We also discuss various complexity aspects of the classification algorithms Xian et al., (2023) and Wang et al., (2023) and relate them to the complexity of MIFPO.
In Section 4.2 we described a data independent discretization of $\Delta_{\mathcal{Y}}$ by binning. While effective for small label sets $\mathcal{Y}$, larger sets would require a different approach. One alternative is to cluster the data instead of binning $\Delta_{\mathcal{Y}}$ itself. Indeed, by choosing cluster centers $\left\{\eta_{i}\right\}_{1}^{M}\subset\Delta_{\mathcal{Y}}$, such that each data point $\left\{f^{*}(x,a)\right\}_{(x,a)\in D}$ is well approximated by one the centers (or just most points are approximated), we can guarantee arbitrarily good approximation of the true Pareto front. The cardinality $M$ of such clustering would depend on the intrinsic dimension of the data, which we typically expect to be lower than the full dimension of $\Delta_{\mathcal{Y}}$, due to the Manifold Hypothesis, Fefferman et al., (2016).
Another possibility is to not use any explicit discretization, and instead to use each data point as a separate bin (equivalently, we use the points themselves as the cluster centers $\left\{\eta_{i}\right\}$). In this case, the complexity scales with the size of the dataset, but not with the dimensions of $\Delta_{\mathcal{Y}}$. The MIFPO construction in Section 4.1 implies that MIFPO in this case would have $O(N^{2})$ variables for a dataset of size $N$. While not applicable for large $N$, this is similar to the complexity of a variety of often used algorithms. Classical example of such complexity is the Spectral Clustering. We also observe that Xian et al., (2023), a recent state of the art fairness classification algorithm mentioned above, also has such complexity.
Indeed, the approach in Xian et al., (2023) involves the computation of a certain transportation plan between data points, and the encoding of such plans also requires $O(N^{2})$ variables. Thus problem sizes occuring in MIFPO would be smaller or equal to those in Xian et al., (2023), despite the fact that MIFPO is solving a considerably more general problem (see Section 2).
Finally, the algorithm in Wang et al., (2023) involves optimization in the space of confusion matrices, with dimensions of size $\left|\mathcal{A}\right|\cdot\left|\mathcal{Y}\right|\times\left|\mathcal{Y}\right|$. As discussed in Section 2, the reduction of the problem to confusion matrices of possible due to special properties of the classification problem and the accuracy loss.
The algorithm in Wang et al., (2023), FairFront, is an iterative algorithm, where each iteration involves solving a certain difference-of-convex (DC) program which is constructed from a full dataset. The class of DC programs is equivalent to that convex-concave programs considered in this paper (see Shen et al., (2016)). In fact, similarly to MIFPO, the algorithm in Wang et al., (2023) also uses DCCP, Shen et al., (2016), although applied to a different problem.
In each iteration, the solution of DC program is then used to add new constraints to a certain main convex program. While it is proved that asymptotically this process converges to the optimal front, there are no bounds on the number of iterations. This may lead to the convex solver crashing due to too many constraints, and in fact we have observed such crashes in our evaluation.
To summarize  [^4] In this Section we have discussed the theoretical complexity aspects of FairFront. Additional details pertaining to the official implementation of Wang et al., (2023) may be found in Section H.2., MIFPO involves solving one convex-concave problem, with size which may be independent of the data size. In contrast, FairFront involves iteratively solving convex-concave problems and a main convex program, where the number of terms in the objective of each convex-concave problem scales with the size of the data, and the number of constraints grows in the convex problem grows with iterations, thus making the iterations progressively harder.
Appendix K Factorization
In this Section we the factorization result, Lemma 3.2. We restate the result for convenience.
Recall that $f^{*}(x,a)$ denotes the Bayes optimal classifier of $Y$, i.e. $f^{*}:\mathcal{X}\times\mathcal{A}\rightarrow\Delta_{\mathcal{Y}}$ is given by
$f^{*}(x,a):=\mathbb{P}\left(Y=\cdot|X=x,A=a\right).$ (117)
We define a new variable $X^{\prime}$, with values in $\Delta_{\mathcal{Y}}$, by $X^{\prime}:=f^{*}(X,A)$.
Lemma K.1 (Factorization).
For any representation $Z$ of $(X,A)$, there is a representation $Z^{\prime}$ of $(X^{\prime},A)$, such that
$\mathbb{E}_{z^{\prime}\sim Z^{\prime}}h(\mathbb{P}\left(Y|Z^{\prime}=z^{\prime}\right))=\mathbb{E}_{z\sim Z}h(\mathbb{P}\left(Y|Z=z\right))\text{ and }D_{TV}(Z^{\prime})\leq D_{TV}(Z).$ (118)
Proof.
The representation $Z^{\prime}$ of $X^{\prime},A$ will be defined as follows: For $\sigma\in\Delta_{\mathcal{Y}},a\in\mathcal{A},z\in\mathcal{Z}$ set:
$\mathbb{P}\left(Z^{\prime}=z|X^{\prime}=\sigma,A=a\right)$ $:=\mathbb{P}\left(Z=z|X^{\prime}=\sigma,A=a\right)$ (119)
$=\mathbb{P}\left(Z=z|f^{*}(X,a)=\sigma,A=a\right)$ (120)
where the second line is the definition and is added for clarity. To see intuition behind this definition note that we have
$\mathbb{P}\left(Z^{\prime}=z|X^{\prime}=\sigma,A=a\right)$ $=\mathbb{P}\left(Z=z|f^{*}(X,a)=\sigma,A=a\right)$ (121)
$=\sum_{x\in\mathcal{X}}\mathbb{P}\left(Z=z|X=x,f^{*}(x,a)=\sigma,A=a\right)\mathbb{P}\left(X=x|f^{*}(x,a)=\sigma,A=a\right)$
$=\sum_{x\in\mathcal{X}}\mathbb{P}\left(Z=z|X=x,A=a\right)\mathbb{P}\left(X=x|f^{*}(x,a)=\sigma,A=a\right)$ (122)
In words, for a fixed $a$, to compute $\mathbb{P}\left(Z^{\prime}=z|X^{\prime}=\sigma,A=a\right)$ we effectively collect all $x$ such that $f^{*}(x,a)=\sigma$ and average all of their representations. Equivalently, all points $x$ with the same $\sigma$ are merged into one point, and their representations summed up according to their relative weight.
We will now show that neither the performance nor the fairness condition change under this operation.
Since $D_{TV}(Z)$ is defined solely in terms of the distributions $\mathbb{P}\left(Z=\cdot|A=a\right)$, to show that $D_{TV}(Z)=D_{TV}(Z^{\prime})$ it is enough to show that
$\mathbb{P}\left(Z=\cdot|A=a\right)=\mathbb{P}\left(Z^{\prime}=\cdot|A=a\right)\hskip 5.69054pt\forall a\in\mathcal{A}.$ (123)
To this end, we have
$\mathbb{P}\left(Z^{\prime}=z|A=a\right)$ $=\sum_{\sigma}\mathbb{P}\left(Z^{\prime}=z|X^{\prime}=\sigma,A=a\right)\mathbb{P}\left(X^{\prime}=\sigma|A=a\right)$ (124)
$=\sum_{\sigma}\mathbb{P}\left(Z=z|X^{\prime}=\sigma,A=a\right)\mathbb{P}\left(X^{\prime}=\sigma|A=a\right)$ (125)
$=\mathbb{P}\left(Z=z|A=a\right).$ (126)
where the transition (124) to (125) is by the definition of $Z^{\prime}$. Thus we have shown that $D_{TV}(Z)=D_{TV}(Z^{\prime})$.
Next, note that the above argument implies also that $\mathbb{P}\left(Z=z\right)=\mathbb{P}\left(Z^{\prime}=z\right)$. Thus, in order to show the performance equality,
$\mathbb{E}_{z^{\prime}\sim Z^{\prime}}h(\mathbb{P}\left(Y|Z^{\prime}=z^{\prime}\right))=\mathbb{E}_{z\sim Z}h(\mathbb{P}\left(Y|Z=z\right)),$ (127)
it is enough to show that $\mathbb{P}\left(Y|Z^{\prime}=z\right)=\mathbb{P}\left(Y|Z=z\right)$ for every $z\in\mathcal{Z}$. Further, again since $\mathbb{P}\left(Z=\cdot\right)=\mathbb{P}\left(Z^{\prime}=\cdot\right)$, we can show that
$\mathbb{P}\left(Y=y,Z^{\prime}=z\right)=\mathbb{P}\left(Y=y,Z=z\right)$ (128)
for all $z\in\mathcal{Z},y\in\mathcal{Y}$. Write
$\mathbb{P}\left(Y=y,Z=z\right)$ (129)
$=\sum_{a}\sum_{x}\mathbb{P}\left(Y=y,X=x,A=a,Z=z\right)$ (130)
$=\sum_{a}\sum_{x}\mathbb{P}\left(Y=y|X=x,A=a,Z=z\right)\mathbb{P}\left(X=x,A=a,Z=z\right)$ (131)
$=\sum_{a}\sum_{x}\mathbb{P}\left(Y=y|X=x,A=a,Z=z\right)\mathbb{P}\left(Z=z|X=x,A=a\right)\mathbb{P}\left(X=x,A=a\right)$ (132)
$=\sum_{a}\sum_{x}\mathbb{P}\left(Y=y|X=x,A=a\right)\mathbb{P}\left(Z=z|X=x,A=a\right)\mathbb{P}\left(X=x,A=a\right)$ (133)
$=\sum_{a}\sum_{\sigma}\sum_{x|f^{*}(x,a)=\sigma}\mathbb{P}\left(Y=y|X=x,A=a\right)\mathbb{P}\left(Z=z|X=x,A=a\right)\mathbb{P}\left(X=x,A=a\right)$ (134)
$=\sum_{a}\sum_{\sigma}\sigma(y)\mathbb{P}\left(X^{\prime}=\sigma,A=a\right)\sum_{x|f^{*}(x,a)=\sigma}\mathbb{P}\left(Z=z|X=x,A=a\right)\frac{\mathbb{P}\left(X=x,A=a\right)}{\mathbb{P}\left(X^{\prime}=\sigma,A=a\right)}$ (135)
$=\sum_{a}\sum_{\sigma}\sigma(y)\mathbb{P}\left(X^{\prime}=\sigma,A=a\right)\sum_{x|f^{*}(x,a)=\sigma}\mathbb{P}\left(Z=z|X=x,A=a\right)\mathbb{P}\left(X=x|X^{\prime}=\sigma,A=a)\right)$ (136)
$=\sum_{a}\sum_{\sigma}\sigma(y)\mathbb{P}\left(X^{\prime}=\sigma,A=a\right)\mathbb{P}\left(Z^{\prime}=z|X^{\prime}=\sigma,A=a\right)$ (137)
$=\sum_{a}\sum_{\sigma}\sigma(y)\mathbb{P}\left(Z^{\prime}=z,X^{\prime}=\sigma,A=a\right)$ (138)
$=\sum_{a}\sum_{\sigma}\mathbb{P}\left(Y=y|X^{\prime}=\sigma,A=a\right)\mathbb{P}\left(Z^{\prime}=z,X^{\prime}=\sigma,A=a\right)$ (139)
$=\mathbb{P}\left(Y=y,Z^{\prime}=z\right).$ (140)
Here, the transition (132)-(133) is due to the independence condition (1). On line (134) we split the sum over $x$ into sum over subsets $\left\{x\hskip 5.69054pt|\hskip 5.69054ptf^{*}(x,a)=\sigma\right\}$ and an outer some over all $\sigma$. The transition (136)-(137) is due to the equality (122)-(121). Finally, for the transition (137)-(138), we have used the fact that
$\sigma(y)=\mathbb{P}\left(Y=y|X^{\prime}=\sigma,A=a\right),$ (141)
which holds by definition of $X^{\prime}$. Similarly to the earlier discussion on merging of $x$ with similar value of $f^{*}$, the above argument proceeded by regrouping the summation over $x$ by the value of $f^{*}(x,a)$. The computation thus showed that this process yields the definition of $Z^{\prime}$. In particular, this regrouping process and equation (141) explain why the space $\Delta_{\mathcal{Y}}$ is special and all representations may be factored through it. This completes the proof of the Lemma. ∎
In the above argument we have used the summation over $\sigma$, i.e. $\sum_{\sigma}\ldots$. This is formally possible when $(X,A)$ has a discreet distribution. The full general case may be obtained simply by replacing the summation by integration and conditioning on $\sigma$ by the general conditional expectation operator.